Hello, I'm really struggling on volume and surface area. I have a word problem I have to solve in a worksheet and I'm having trouble. Anyways, the problem is stated exactly as follows: "Prism A and prism B are similar with a ratio of 2:3. If the volume of prism A is 36 cubic units, what is the volume of prism B?
If the surface area of prism A is 90 square units, what is the surface area of prism B?" Thanks.

Answer by scott8148(6628)

an example may be helpful

suppose we have two similar cubes (yes, ALL cubes are similar)
one with an edge length of x
and the other with edge length 2x
the ratio of the cubes is x:2x or 1:2

the surface area of the smaller cube is 6*x^2 or 6x^2
the surface area of the larger cube is 6*(2x)^2 or 24x^2
the ratio of the surface areas is 6x^2:24x^2 or 6:24 or 1:4 or (1:2)^2

the volume of the smaller cube is x^3
the volume of the larger cube is (2x)^3 or 8x^3
the ratio of the volumes is x^3:8x^3 or 1:8 or (1:2)^3

the ratio of the surface areas is the SQUARE of the proportion ratio
and the ratio of the volumes is the CUBE of the proportion ratio



for the prism ___ A:B = 2:3

(A:B)^3 = (2:3)^3 = 8:27

let x = volume B
___ 8/27 = 36/x
___ "cross" multiplying ___ 8x = 27*36
___ dividing by 8 ___ x = 121.5


(A:B)^2 = (2:3)^2 = 4:9

let y = surface area B
___ 4/9 = 90/x
___ "cross" multiplying ___ 4x = 810
___ dividing by 4 ___ x = 202.5